OpenStudy (anonymous):

Solve the logarithmic equation. log(x + 9) - logx = 3

OpenStudy (anonymous):

x+9/x=1000 1+9/x=1000 x=1/111

OpenStudy (anonymous):

\[\log(x+9)-logx=\log_{10}\frac{x+9}{x}=3\]

OpenStudy (anonymous):

\[\frac{x+9}x=10^3\]\[1+\frac9x=1000\]\[\frac9x=999\]\[x=\frac{9}{999}=\frac1{111}\]

OpenStudy (anonymous):

if base of logarithm is 10, log(x + 9) - logx = 3 log(1 +9/x) = 3 1+9/x =1000 9/x=999 x=1/111